# On the number of linear uniform hypergraphs with girth constraint

### v1

@@ Abstract
For an integer $r\geqslant 3$, a hypergraph on vertex set $[n]$ is $r$-uniform if each edge is a set of $r$ vertices, and is said to be linear if every two distinct edges share at most one vertex. Given a family $\mathcal{H}$ of linear $r$-uniform hypergraphs, let $\textmd{Forb}_{r}^{L}(n,\mathcal{H})$ be the set of linear $r$-uniform hypergraphs on vertex set $[n]$, which do not contain any member from $\mathcal{H}$ as a subgraph. An $r$-uniform linear cycle of length $\ell$, denoted by $C_{\ell}^{r}$, is a linear $r$-uniform hypergraph on $(r-1)\ell$ vertices whose edges can be ordered as $\boldsymbol{e}_{1},\ldots,\boldsymbol{e}_{\ell}$ such that $|\boldsymbol{e}_{i}\cap\boldsymbol{e}_{j}|=1$ if $j=i\pm 1$ (indices taken modulo $\ell$) and $|\boldsymbol{e}_{i}\cap\boldsymbol{e}_{j}|=0$ otherwise. The girth of a linear $r$-uniform hypergraph is the smallest integer $\ell$ such that it contains a $C_{\ell}^{r}$. Let $\textmd{Forb}_{L}(n,r,\ell)=\textmd{Forb}_{r}^{L}(n,\mathcal{H})$ when $\mathcal{H}=\{C_{i}^{r}:\,3\leqslant i\leqslant\ell\}$, that is, $\textmd{Forb}_{L}(n,r,\ell)$ is the set of all linear $r$-uniform hypergraphs on $[n]$ with girth larger than $\ell$. For integers $r\geqslant 3$ and $\ell\geqslant 4$, Balogh and Li [On the number of linear hypergraphs of large girth, J. Graph Theory, 93(1) (2020), 113-141] showed that $|\textmd{Forb}_{L}(n,r,\ell)|=2^{O(n^{1+1/\lfloor\ell/2\rfloor})}$ based on the graph container method, while its sharpness remains open. In this paper, we prove that $|\textmd{Forb}_{L}(n,r,\ell)|>2^{n^{1+1/(\ell-1)-O(\log\log n/\log n)}}$ by analyzing the random greedy high girth linear $r$-uniform hypergraph process. It partially generalizes some known results on linear Turán number of linear cycles in higher uniformities.
Keywords: linear hypergraphs, linear Turán number, linear cycles, girth, random greedy graph process.
@@ 1 Introduction
$2^{\textmd{ex}_{r}^{L}(n,\mathcal{H})}\leqslant|\textmd{Forb}_{r}^{L}(n,\mathcal{H})|\leqslant\sum_{i\leqslant\textmd{ex}_{r}^{L}(n,\mathcal{H})}\binom{\binom{n}{r}}{i}\leqslant 2n^{r\cdot\textmd{ex}_{r}^{L}(n,\mathcal{H})}.$
There are several different notions of hypergraph cycles. For an integer $\ell\geqslant 3$, a Berge cycle $B_{\ell}^{r}$ of length $\ell$ is an alternating sequence of distinct vertices and distinct edges of the form ${v}_{1},\boldsymbol{e}_{1},{v}_{2},\boldsymbol{e}_{2},\ldots,{v}_{\ell},\boldsymbol{e}_{\ell}$, where ${v}_{i},{v}_{i+1}\in\boldsymbol{e}_{i}$ for each $i\in[\ell-1]$ and ${v}_{\ell},{v}_{1}\in\boldsymbol{e}_{\ell}$. This definition of a hypergraph cycle is the classical definition due to Berge. Note that forbidding a Berge cycle $B_{\ell}^{r}$ actually forbids a family of hypergraphs. Another notion of hypergraph cycles that has been actively investigated is that of a linear cycle, which is unique up to isomorphism. An $r$-uniform linear cycle $C_{\ell}^{r}$ of length $\ell$ is an alternating sequence ${v}_{1},\boldsymbol{e}_{1},{v}_{2},\boldsymbol{e}_{2},\ldots,{v}_{\ell},\boldsymbol{e}_{\ell}$ of distinct vertices and distinct edges such that $\boldsymbol{e}_{i}\cap\boldsymbol{e}_{i+1}=\{{v}_{i+1}\}$ for each $i\in[\ell-1]$, $\boldsymbol{e}_{1}\cap\boldsymbol{e}_{\ell}=\{{v}_{1}\}$ and $\boldsymbol{e}_{i}\cap\boldsymbol{e}_{j}=\emptyset$ if $1<|j-i|<\ell-1$. The girth of a linear $r$-graph is to be the smallest integer $\ell$ such that it contains a $C_{\ell}^{r}$. If no such $\ell$ exists, the girth of the linear $r$-graph is defined to be infinite.
@@
where $\epsilon=0$ if $k$ is odd and $\epsilon=1$ if $k$ is even. As a special case of the linear hypergraph extension of Kővári–Sós–Turán’s theorem in [14], Gao and Chang showed that $\textmd{ex}_{3}^{L}(n,C_{4}^{3})=\frac{1}{6}n^{3/2}+O(n)$, and Gao et al. in [15] further proved that $\textmd{ex}_{3}^{L}(n,C_{5}^{3})=\frac{1}{3\sqrt{3}}n^{3/2}+O(n)$. For the lower bound of $\textmd{ex}_{r}^{L}(n,C_{\ell}^{r})$, the best known lower bound for $r=3$ is
$\textmd{ex}_{3}^{L}\bigl(n,C_{\ell}^{3}\bigr)=\Omega(n^{1+\frac{1}{\ell-1}})$
@@
that was shown in [9, 11] due to Verstraëte, by taking a random subgraph of a Steiner triple system. Using generalized Sidon sets [11], it is conceivable that one can obtain a similar (or better) constructive lower bound of $\textmd{ex}_{3}^{L}(n,C_{\ell}^{3})$ (and maybe also for all $r\geqslant 3$), which is an area worth some exploration. The lower bound on the linear Turán number of linear cycles is still far from what is conjectured, especially in higher uniformities $r\geqslant 4$. Sometimes to explore the lower bound of the corresponding Turán problem is more difficult than the upper one.
@@
Following the same logic with the Turán problems of cycles, for integers $r\geqslant 3$ and $\ell\geqslant 4$, Balogh and Li [2] conjectured that $|\textmd{Forb}_{r}^{L}(n,C_{\ell}^{r})|=2^{\Theta(n^{1+1/\lfloor\ell/2\rfloor})}$. They confirmed the conjecture in the case of $\ell=4$. For $\ell>4$, they provided a result on the girth version, that is, there exists a constant $c=c(r,\ell)$ such that the number of linear $r$-graphs with girth larger than $\ell$ is at most $2^{c\cdot n^{1+1/\lfloor\ell/2\rfloor}}$. Define $\textmd{ex}_{L}(n,r,\ell)$ and $\textmd{Forb}_{L}(n,r,\ell)$ as $\textmd{ex}_{r}^{L}(n,\mathcal{H})$ and $\textmd{Forb}_{r}^{L}(n,\mathcal{H})$ of $\mathcal{H}=\{C_{i}^{r}:\,3\leqslant i\leqslant\ell\}$, respectively. Hence, it implies that $|\textmd{Forb}_{L}(n,r,\ell)|=2^{O(n^{1+1/\lfloor\ell/2\rfloor})}$, while its lower bound remains open. The idea of the proof in [2] is to reduce the hypergraph enumeration problems to some graph enumeration problems, and then the graph container method is implemented in it. Balogh, Narayanan and Skokan [1] indeed made use of the hypergraph container method and provided a balanced supersaturation theorem for linear cycles to resolve a conjecture on the enumeration of $r$-uniform hypergraphs with a forbidden $C_{\ell}^{r}$.
@@
In order to solve a conjecture of Erdős about the existence of sparse Steiner triple systems, Glock, Kühn, Lo and Osthus [16], and Bohman and Warnke [7], studied a constrained triple process along similar lines, which can be seen as a generalization of the random greedy triangle removal process [6], to iteratively build a sparse partial Steiner triple system, that is, adding the sparseness constraint does not affect the evolution of the process significantly. To be precise, a Steiner triple system is called $k$-sparse if it does not contain any $(i+2,i)$-configurations for integers $2\leqslant i\leqslant k$, where a $(i+2,i)$-configuration is a set of $i$ triples which span at most $i+2$ vertices. Erdős made his conjecture in Steiner triple systems. Some further problems to higher uniformities have been proposed and considered in [13, 16].
@@
The random greedy $r$-clique removal process starts with a complete graph on vertex set $[n]$, denoted as $\mathbb{G}(0)$, and $\mathbb{G}(i+1)$ is the remaining graph from $\mathbb{G}(i)$ by selecting one $r$-clique uniformly at random out of all $r$-cliques in $\mathbb{G}(i)$ and deleting all its edges from the edge set E(i){E}(i) of $\mathbb{G}(i)$. The process terminates once the remaining graph contains no $r$-cliques, and along the way produces a linear $r$-graph, where it starts with the empty $r$-graph on vertex set $[n]$, iteratively adding the vertex set of each chosen $r$-clique as a new edge in the hypergraph. Let $M=\min\{i\,:\mathbb{G}(i)\ {\mbox{is\ }r{\mbox{-clique free}}}\}$, and E(M){E}(M) be the set of edges left unsaturated by the produced $r$-cliques. In particular, this problem attracted much attention when $r=3$ and it is also called the random greedy triangle removal process. Fewer results are known on $r\geqslant 4$. Bennett and Bohman [4] considered the random greedy hypergraph matching process, which can be seen as a generalization of the $r$-clique removal process. Tian et al. [23] directly discussed the structure of the random greedy $r$-clique removal process to a natural barrier, using a heuristic assumption to find the trajectories of an ensemble of random variables when the process evolves.
@@
Inspired by the above works, for integers $r\geqslant 3$ and $\ell\geqslant 3$, we consider the constrained random greedy $r$-clique removal process so that it does not just produce a linear $r$-graph but with girth larger than $\ell$, and the corresponding process is called the random greedy high girth $r$-clique removal process or high girth linear $r$-uniform hypergraph process. Compared with their analysis in [16, 7] discussing sparse Steiner triple systems, the challenge arising in the analysis of random greedy high girth $r$-clique removal process is that each individual forbidden configuration $C_{i}^{r}$ with $3\leqslant i\leqslant\ell$ should be considered. It is more sensitive and stricter to establish the trajectories and error functions of an ensemble of random variables which we track in this constrained process with a higher uniformity, such that the one-step changes, trend hypotheses and boundedness hypotheses of these tracked variables can support the proof about the concentration of all these variables to the end. Finally, we show the following theorem.
@@
For every pair of fixed integers $r,\ell\geqslant 3$, consider the random greedy high girth $r$-clique removal process on vertex set $[n]$. Let $M$ be the number of edges in the generated linear $r$-graph with girth larger than $\ell$ when the process terminates. With high probability, there exists some positive constant $\lambda=\lambda(\ell)$ such that
@@
In the following proof of Theorem 1.1, we will show the term $O({\log\log n}/{\log n})$ is necessary in the exponent of the lower bound of $M$. Furthermore, it is impossible to apply this approach to obtain a better lower bound than $\Omega(n^{1+\frac{1}{\ell-1}})$. As a corollary of Theorem 1.1, it is natural to obtain lower bounds of $\textmd{ex}_{L}(n,r,\ell)$ and $|\textmd{Forb}_{L}(n,r,\ell)|$ when $\ell\geqslant 3$.
@@
For every pair of fixed integers $r,\ell\geqslant 3$, there exists $\lambda=\lambda(\ell)>0$ such that as $n\rightarrow\infty$,
$\textmd{ex}_{L}(n,r,\ell)>n^{1+\frac{1}{\ell-1}-\frac{\lambda\log\log n}{\log n}}\quad\text{and}\quad|\textmd{Forb}_{L}(n,r,\ell)|>2^{n^{1+\frac{1}{\ell-1}-\frac{\lambda\log\log n}{\log n}}}.$
The remainder of the paper is structured as follows. Notation and auxiliary results used throughout the paper are presented in Section 2. We define the random greedy high girth $r$-clique removal process in Section 3, introducing some key random variables of the process that we wish to track, estimating the corresponding expected trajectory and choosing error function for each tracked variable. We formally prove the concentration of all these variables in Section 4, where the required one-step changes, trend hypotheses and boundedness hypotheses for these tracked variables are analyzed in higher uniformities. Two important claims to bound the overcounting and boundedness parameters are discussed in detail in the final section.

### v2

# On the number of linear uniform hypergraphs with linear girth constraint

@@ Abstract
For an integer $r\geqslant 3$, a hypergraph on vertex set $[n]$ is $r$-uniform if each edge is a set of $r$ vertices, and is said to be linear if every two distinct edges share at most one vertex. Given a family $\mathcal{H}$ of linear $r$-uniform hypergraphs, let $\textmd{Forb}_{r}^{L}(n,\mathcal{H})$ be the set of linear $r$-uniform hypergraphs on vertex set $[n]$, which does not contain any member from $\mathcal{H}$ as a subgraph. An $r$-uniform linear cycle of length $\ell$, denoted by $C_{\ell}^{r}$, is a linear $r$-uniform hypergraph on $(r-1)\ell$ vertices whose edges can be ordered as $\boldsymbol{e}_{1},\ldots,\boldsymbol{e}_{\ell}$ such that $|\boldsymbol{e}_{i}\cap\boldsymbol{e}_{j}|=1$ if $j=i\pm 1$ (indices taken modulo $\ell$) and $|\boldsymbol{e}_{i}\cap\boldsymbol{e}_{j}|=0$ otherwise. The linear girth of a linear $r$-uniform hypergraph is the smallest integer $\ell$ such that it contains a $C_{\ell}^{r}$. Let $\textmd{Forb}_{L}(n,r,\ell)=\textmd{Forb}_{r}^{L}(n,\mathcal{H})$ when $\mathcal{H}=\{C_{i}^{r}:\,3\leqslant i\leqslant\ell\}$, that is, $\textmd{Forb}_{L}(n,r,\ell)$ is the set of all linear $r$-uniform hypergraphs on $[n]$ with linear girth greater than $\ell$. For integers $r\geqslant 3$ and $\ell\geqslant 4$, Balogh and Li [On the number of linear hypergraphs of large girth, J. Graph Theory, 93(1) (2020), 113-141] showed that $|\textmd{Forb}_{L}(n,r,\ell)|=2^{O(n^{1+1/\lfloor\ell/2\rfloor})}$ based on the graph container method. It is natural to obtain $|\textmd{Forb}_{L}(n,r,\ell)|\geqslant 2^{c\cdot n^{1+1/\ell}}$ for some constant $c$ by probabilistic deletion method. Combined with the known results that $|\textmd{Forb}_{L}(n,r,3)|=2^{o(n^{2})}$ and $|\textmd{Forb}_{L}(n,3,4)|=2^{\Theta(n^{3/2})}$, by analyzing the random greedy high linear girth linear $r$-uniform hypergraph process, we show $|\textmd{Forb}_{L}(n,r,\ell)|\geqslant 2^{n^{1+1/(\ell-1)-O(\log\log n/\log n)}}$ for every pair of fixed integers $r,\ell\geqslant 4$, or $r=3$ and $\ell\geqslant 5$.
Keywords: linear hypergraphs, linear Turán number, linear cycles, linear girth, random greedy graph process.
@@ 1 Introduction
$2^{\textmd{ex}_{r}^{L}(n,\mathcal{H})}\leqslant|\textmd{Forb}_{r}^{L}(n,\mathcal{H})|\leqslant\sum_{i\leqslant\textmd{ex}_{r}^{L}(n,\mathcal{H})}\binom{\binom{n}{r}}{i}\leqslant 2n^{r\cdot\textmd{ex}_{r}^{L}(n,\mathcal{H})}.$ (1.1)
There are several different notions of hypergraph cycles. For an integer $\ell\geqslant 3$, a Berge cycle $B_{\ell}^{r}$ of length $\ell$ is an alternating sequence of distinct vertices and distinct edges of the form ${v}_{1},\boldsymbol{e}_{1},{v}_{2},\boldsymbol{e}_{2},\ldots,{v}_{\ell},\boldsymbol{e}_{\ell}$, where ${v}_{i},{v}_{i+1}\in\boldsymbol{e}_{i}$ for each $i\in[\ell-1]$ and ${v}_{\ell},{v}_{1}\in\boldsymbol{e}_{\ell}$. This definition of a hypergraph cycle is the classical definition due to Berge. Note that forbidding a Berge cycle $B_{\ell}^{r}$ actually forbids a family of hypergraphs. Another notion of hypergraph cycles that has been actively investigated is that of a linear cycle, which is unique up to isomorphism. An $r$-uniform linear cycle $C_{\ell}^{r}$ of length $\ell$ is an alternating sequence ${v}_{1},\boldsymbol{e}_{1},{v}_{2},\boldsymbol{e}_{2},\ldots,{v}_{\ell},\boldsymbol{e}_{\ell}$ of distinct vertices and distinct edges such that $\boldsymbol{e}_{i}\cap\boldsymbol{e}_{i+1}=\{{v}_{i+1}\}$ for each $i\in[\ell-1]$, $\boldsymbol{e}_{1}\cap\boldsymbol{e}_{\ell}=\{{v}_{1}\}$ and $\boldsymbol{e}_{i}\cap\boldsymbol{e}_{j}=\emptyset$ if $1<|j-i|<\ell-1$. The linear girth of a linear $r$-graph is the smallest integer $\ell$ such that it contains a $C_{\ell}^{r}$. If no such $\ell$ exists, the linear girth of the linear $r$-graph is defined to be infinite.
@@
where $\epsilon=0$ if $k$ is odd and $\epsilon=1$ if $k$ is even. As a special case of the linear hypergraph extension of Kővári–Sós–Turán’s theorem in [17], Gao and Chang showed that $\textmd{ex}_{3}^{L}(n,C_{4}^{3})=\frac{1}{6}n^{3/2}+O(n)$, and Gao et al. in [18] further proved that $\textmd{ex}_{3}^{L}(n,C_{5}^{3})=\frac{1}{3\sqrt{3}}n^{3/2}+O(n)$. For the lower bound of $\textmd{ex}_{r}^{L}(n,C_{\ell}^{r})$, the best known lower bound for $r=3$ is $\textmd{ex}_{3}^{L}\bigl(n,C_{\ell}^{3}\bigr)=\Omega(n^{1+\frac{1}{\ell-1}})$ that was mentioned in [10, 14] due to Verstraëte, by taking a random subgraph of a Steiner triple system. The lower bound on the linear Turán number of linear cycles is still far from what is conjectured, especially in a higher uniformity $r\geqslant 4$, which is an area worth some exploration [14]. Sometimes to explore the lower bound of the corresponding Turán problem is more difficult than the upper one.
@@
Following the same logic with the Turán problems of cycles, for integers $r\geqslant 3$ and $\ell\geqslant 4$, Balogh and Li [2] conjectured that $|\textmd{Forb}_{r}^{L}(n,C_{\ell}^{r})|=2^{\Theta(n^{1+1/\lfloor\ell/2\rfloor})}$. For $\ell=3$, the work of Erdős, Frankl and Rödl [13] could be extended to show that $|\textmd{Forb}_{r}^{L}(n,C_{3}^{r})|=2^{o(n^{2})}$. Balogh and Li [2] confirmed the conjecture in the case of $\ell=4$. In fact, they proved that for every $r\geqslant 3$ there exists $C=C(r)>0$ such that $|\textmd{Forb}_{r}^{L}(n,C_{4}^{r})|\leqslant 2^{C\cdot n^{3/2}}$. The upper bound for $C_{4}^{3}$ is sharp in order of magnitude given by $\textmd{ex}_{r}^{L}(n,C_{4}^{3})=\Theta(n^{3/2})$ and (1.1). In general, for $r\geqslant 3$ and $\ell\geqslant 4$, they provided a result on the girth version, that is, there exists a constant $c=c(r,\ell)$ such that the number of linear $r$-graphs with linear girth greater than $\ell$ is at most $2^{c\cdot n^{1+1/\lfloor\ell/2\rfloor}}$. Define $\textmd{ex}_{L}(n,r,\ell)$ and $\textmd{Forb}_{L}(n,r,\ell)$ as $\textmd{ex}_{r}^{L}(n,\mathcal{H})$ and $\textmd{Forb}_{r}^{L}(n,\mathcal{H})$ of $\mathcal{H}=\{C_{i}^{r}:\,3\leqslant i\leqslant\ell\}$, respectively. Hence, it implies that $|\textmd{Forb}_{L}(n,r,\ell)|=2^{O(n^{1+1/\lfloor\ell/2\rfloor})}$. They stated that the upper bounds are possibly sharp, while they are not able to confirm it now. The idea of the proof in [2] is to reduce the hypergraph enumeration problems to some graph enumeration problems, and then the graph container method is implemented in them. Balogh, Narayanan and Skokan [1] made use of the hypergraph container method and provided a balanced supersaturation theorem for linear cycles to resolve a conjecture on the enumeration of $r$-uniform hypergraphs with a forbidden $C_{\ell}^{r}$.
@@
In order to solve a conjecture of Erdős about the existence of sparse Steiner triple systems, Glock, Kühn, Lo and Osthus [19], and Bohman and Warnke [8], studied a constrained triple process along similar lines, which can be seen as a generalization of the random greedy triangle removal process [7], to iteratively build a sparse partial Steiner triple system, that is, adding the sparseness constraint does not affect the evolution of the process significantly. To be precise, a Steiner triple system is called $\ell$-sparse, or girth greater than $\ell$, if it does not contain any $(i+2,i)$-configurations for integers $2\leqslant i\leqslant\ell$, where a $(i+2,i)$-configuration is a set of $i$ triples which span at most $i+2$ vertices. Erdős made his conjecture in Steiner triple systems. Kwan, Sah, Sawhney and Simkin [26] proved Erdős’s conjecture with proof built upon the approximate results of [8, 19] and the iterative absorption method, combined with many new ideas. Erdős’s conjecture for Steiner systems in higher uniformities was conjectured in several papers, such as  [16, 19, 24]. Delcourt and Postle [11] and independently Glock, Joos, Kim, Kühn and Lichev [20] settled approximate versions of the conjecture made in [16, 19, 24]. Both approaches offer different benefits, and each has led to substantial subsequent work. In a very recent breakthrough, Delcourt and Postle [12] proved Erdős’s conjecture for general Steiner systems.
@@
The random greedy $r$-clique removal process starts with a complete graph on vertex set $[n]$, denoted as $\mathbb{G}(0)$, and $\mathbb{G}(i+1)$ is the remaining graph from $\mathbb{G}(i)$ by selecting one $r$-clique uniformly at random out of all $r$-cliques in $\mathbb{G}(i)$ and deleting all its edges from the edge set E(i){E}(i) of $\mathbb{G}(i)$. The process terminates once the remaining graph contains no $r$-cliques, and along the way produces a linear $r$-graph, where it starts with the empty $r$-graph on vertex set $[n]$, iteratively adding the vertex set of each chosen $r$-clique as a new edge in the hypergraph. Let $M=\min\{i\,:\mathbb{G}(i)\ {\mbox{is\ }r{\mbox{-clique free}}}\}$, and E(M){E}(M) be the set of edges left unsaturated by the produced $r$-cliques, which are related via
$|E(M)|=\binom{n}{2}-\binom{r}{2}M.$
In particular, this problem attracted much attention when $r=3$ and it is also called the random greedy triangle removal process [7]. Fewer results are known on $r\geqslant 4$. Bennett and Bohman [4] considered the random greedy hypergraph matching process, which can be seen as a generalization of the $r$-clique removal process. A special case of the conjecture proposed in [4] is w.h.p. $|E(M)|=n^{2-\frac{2}{r+1}+o(1)}$. Their upper bound still has not the correct order of magnitude. Without the appropriate substructures, it seems impossible to rely on self-correcting behavior to determine the order of $|E(M)|$. Tian et al. [31] directly discussed the structure of the random greedy $r$-clique removal process to the same natural barrier with the one in [4], that is, $|E(M)|\leqslant n^{2-\frac{1}{r(r-1)-2}+o(1)}$. This upper bound on $|E(M)|$ equals $n^{\frac{7}{4}+o(1)}$ when $k=3$, coinciding with an upper bound obtained in [21, 6] for random triangle packing. Joos and Kühn [23] finally solved this conjecture in a significantly strong form. Therefore, the maximum number of edges in a linear $r$-graph on $n$ vertices lies between $\binom{n}{2}/\binom{r}{2}-O(n^{2-\frac{2}{r+1}+o(1)})$ and $\binom{n}{2}/\binom{r}{2}$.
By classical probabilistic deletion method, for every pair of integers $r,\ell\geqslant 3$, it is natural to prove that there exists a linear $r$-graph on vertex set $[n]$ that has at least $c\cdot n^{1+\frac{1}{\ell}}$ edges for some constant $c$, such as $c=\frac{1}{2r(r-1)}$, and the linear girth greater than $\ell$, which implies that $\textmd{ex}_{L}(n,r,\ell)\geqslant c\cdot n^{1+\frac{1}{\ell}}$ and $|\textmd{Forb}_{L}(n,r,\ell)|\geqslant 2^{c\cdot n^{1+\frac{1}{\ell}}}$. In fact, it also shows that $\textmd{ex}_{r}^{L}(n,C_{\ell}^{r})\geqslant c\cdot n^{1+\frac{1}{\ell}}$ and $|\textmd{Forb}_{r}^{L}(n,C_{\ell}^{r})|\geqslant 2^{c\cdot n^{1+\frac{1}{\ell}}}$. We choose a random subgraph of a linear $r$-graph, where each edge is generated independently with probability $p=n^{-\alpha}$ with $\alpha=1-\frac{1}{\ell}$. The expected number of edges and the expected number of linear cycles of length no more than $\ell$ are approximately $\frac{1}{r(r-1)}n^{2-\alpha}$ and $\Theta(n^{\ell(1-\alpha)})$, respectively. We delete one edge from each of these linear cycles to obtain one new linear $r$-graph such that the linear girth of it is greater than $\ell$, and the number of edges in it is at least $\frac{1}{2r(r-1)}n^{1+\frac{1}{\ell}}$, while we can not have the number of edges is $o(n^{1+\frac{1}{\ell-1}})$ by taking $p=o(n^{-\frac{\ell-2}{\ell-1}})$ in this way.
@@
Inspired by the ideas in addressing Erdős conjecture for Steiner systems, especially the ones in [8, 19], we considered the constrained random greedy $r$-clique removal process to enumerate $|\textmd{Forb}_{L}(n,r,\ell)|$. As stated in [23], such random processes are easy to formulate, in many cases however, a precise analysis is challenging. The central questions often concern structural properties that typically hold for the objects generated at termination. For integers $r\geqslant 3$ and $\ell\geqslant 3$, the constrained process does not just produce a linear $r$-graph but with linear girth greater than $\ell$, and the corresponding process is called the random greedy high linear-girth $r$-clique removal process or high linear-girth linear $r$-uniform hypergraph process. Compared with their analysis in [19, 8] discussing sparse Steiner triple systems, the challenge arising in the analysis of random greedy high linear-girth $r$-clique removal process is that when each individual forbidden configuration $C_{i}^{r}$ with $3\leqslant i\leqslant\ell$ is considered, it is more sensitive and stricter to establish the trajectories and error functions of an ensemble of random variables which we track in a higher uniformity, such that the one-step changes, trend hypotheses and boundedness hypotheses of these tracked variables can support the proof about the concentration of all these variables to the end. Finally, we show the following theorem.
@@
For every pair of fixed integers $r,\ell\geqslant 3$, consider the random greedy high linear-girth $r$-clique removal process on vertex set $[n]$. Let $M$ be the number of edges in the generated linear $r$-graph with linear girth greater than $\ell$ when the process terminates. With high probability, there exists some positive constant $\lambda=\lambda(\ell)$ such that
@@
According to the proof of Theorem 1.1 in Section 4, we will show the term $O({\log\log n}/{\log n})$ can not be removed from the exponent of the lower bound of $M$ and it is sufficient in our proof to show the one-step changes, trend hypotheses and boundedness hypotheses of these tracked variables satisfy the conditions in Freedman’s martingale concentration inequality. Furthermore, we will show that applying the random greedy process analysis to obtain a better lower bound of $M$ than $\Omega(n^{1+\frac{1}{\ell-1}})$ is impossible. We believe removing the term $O({\log\log n}/{\log n})$ in Theorem 1.1 entails some new substructures and analysis.
Combined with the known results that $|\textmd{Forb}_{L}(n,r,3)|=2^{o(n^{2})}$ and $|\textmd{Forb}_{L}(n,3,4)|=2^{\Theta(n^{3/2})}$ extended from [2, 13], in this paper, as a corollary of Theorem 1.1, it is natural to obtain that
@@
For every pair of fixed integer $r,\ell\geqslant 4$, or $r=3$ and $\ell\geqslant 5$, there exists $\lambda=\lambda(\ell)>0$ such that as $n\rightarrow\infty$,
$\textmd{ex}_{L}(n,r,\ell)\geqslant n^{1+\frac{1}{\ell-1}-\frac{\lambda\log\log n}{\log n}}\quad\text{and}\quad|\textmd{Forb}_{L}(n,r,\ell)|\geqslant 2^{n^{1+\frac{1}{\ell-1}-\frac{\lambda\log\log n}{\log n}}}.$
The remainder of the paper is structured as follows. Notation and auxiliary results used throughout the paper are presented in Section 2. We define the random greedy high linear-girth $r$-clique removal process in Section 3, introducing some key random variables of the process that we wish to track, estimating the corresponding expected trajectory and choosing error function for each tracked variable. We formally prove the concentration of all these variables in Section 4, where the required one-step changes, trend hypotheses and boundedness hypotheses for these tracked variables are analyzed in higher uniformities. Two important claims to bound the overcounting and boundedness parameters are discussed in detail in the final section.